# ALERT Open Dataset and Input-Size-Agnostic Vision Transformer for Driver Activity Recognition using IR-UWB

### v1

@@ VI-C1 Analysis for manipulation of data size
As shown in Fig. 10(a), we observe that accuracy improves as the observation window increases. Specifically, window sizes of 1 and 2 seconds do not provide sufficient performance due to the limited time available to accurately identify activities. In contrast, window sizes of 5 and 10 seconds offer adequate time to enhance performance. However, longer observation windows also lead to increased inference time and higher computational costs. Therefore, it is essential to consider a balance between observation window size, system responsiveness, and practical computational efficiency.
In Fig. 10(b), we compare the total range bins (about 9 m) with a cropped range bins (<< 2.7 m) to examine the effects of long-delayed multi-path signals. The ability to effectively extract features from long-delayed multi-path signals varies depending on the structure of the algorithm. However, except for ResNet, all algorithms perform better with cropping applied. Although multipath signals may occasionally carry useful information about the surrounding environment—such as the presence of passengers or the type of vehicle—they generally act as noise for most algorithms. Consequently, applying cropping is beneficial for most algorithms.
Additionally, we analyze lower (<< 375 MHz), higher (375–750 MHz), and total frequency areas to determine which segments best capture the movement characteristics associated with various activities. In frequency domain, distinct results emerge for each algorithm, with the ability of feature extraction varying depending on the algorithm used. Consequently, these results should guide the selection of an appropriate algorithm.
@@ VI-C2 Analysis for the number of shots
We present the impact of varying the number of shots on a few-shot adaptation. To obtain a few shot samples, the driver has to capture the activities, which can be burdensome; however, this occurs only once when the system is initially operated. For the few-shot adaptation phase, we utilize various number of shot samples and 10 epochs adaptation, which are enough to fit the model to a driver. As shown in Fig. 10(c), the accuracy increases when all models’ number of shots increases. We observe that even with the addition of only about 5 shots, the accuracy gain is significant. In particular, when applying 30 shots to ISA-ViT, it exhibits a maximum performance of 91.75%.
@@ VI-C4 Analysis for the range-Doppler data domain
As shown in Fig. 2, the ALERT dataset manipulates UWB radar sensing data into range-time and frequency-time domains. One of the key contributions of the ALERT dataset is that users can flexibly manipulate the data according to their specific needs. To demonstrate this data flexibility, we conduct an experiment using range-Doppler domain data. The range-Doppler format is used in datasets such as RaDA, and the ALERT dataset also supports this format. Users can easily perform this conversion using our provided script (ALERT_makeDataset.py) available in the public repository (https://github.com/ALERTdataset/ALERT.git).
As shown in Fig. 12, CNN- and RNN-based methods using range-Doppler inputs do not outperform their counterparts using domain fusion inputs shown in Fig. 9(a). In contrast, ViT-based models, including DeiT and ViT, achieve better performance when using range-Doppler data compared to domain fusion inputs.
This difference can be explained by how the models handle the input data. CNNs and RNNs can directly process UWB radar inputs without resizing. In contrast, DeiT and ViT require resizing the input to $\times 224$, causing a significant reduction in resolution. During domain fusion, the time resolution is cut by more than half in both the range-time and frequency-time domains. Similarly, converting to the range-Doppler domain reduces the Doppler resolution by about half. However, we observe that the loss of time resolution has a more critical impact on these models than the loss of Doppler resolution.
Our proposed model, ISA-ViT, is built on a ViT backbone but incorporates a resizing strategy that preserves the original information without loss. This design enables ISA-ViT to achieve better performance compared to other models when using input from the Doppler range domain. Furthermore, ISA-ViT employs a complementary fusion technique that effectively integrates range and frequency information. As demonstrated in Fig. 9(a), this approach leads to improved performance compared to using range-Doppler data alone.
@@ VI-E1 Overall performance
We refer to the classification accuracy (76.28%), F1 score in Table VII, and confusion matrix in Fig. 14(b) to analyze the model’s accuracy from both safety and service perspectives. The classification accuracy reflects the model’s ability to categorize all activities within our proposed framework. The precision and recall of each activity indicate the safety and service factors. If the recall for Drive is low, the system may trigger false alarms, potentially interrupting the driving experience and reducing overall service quality. While false alarms may be frustrating from a service perspective, they do not pose significant safety risks. However, the precision for Drive is directly linked to safety. Low precision could indicate that other potentially dangerous activities, such as using a smartphone, smoking, nodding are incorrectly classified as driving. Thus, we should refer to classification accuracy and F1 score together to analyze the accuracy in perspectives of safety and service.
@@
TABLE VII: Precision, recall, and F1 score for each label
@@ VI-E2 Impact on resizing method
To evaluate ISA-ViT, we conduct experiments with several resizing methods for the ALERT and RaDA datasets. We have already verified that maintaining the $\times14$ sequence of pre-trained PEVs yields better performance than manipulating them in Section V. Therefore, we only compare methods that maintain pre-trained PEVs.
@@ VI-E3 Impact on domain fusion
TABLE VIII: Comparison of precision, recall, and F1 score before/after domain fusion
@@
Moreover, Fig. 16 illustrates the accuracy impact of the trained adjusting factor $\beta$ for frequency features. Before training the adjusting factor $\beta$, we achieve a performance of 74.02% with range and frequency features at an equal ratio. However, after jointly training $\beta$ and applying it to domain fusion, we observe an accuracy of 76.28%. This confirms that it is crucial not merely to combine range and frequency features, but to balance them appropriately for optimal fusion.
The detailed performance enhancements are presented in Table VIII, which shows the precision, recall, and F1 score for each activity before and after applying domain fusion. For all activities, the F1 scores improve after domain fusion is applied. These improvements demonstrate that by combining information from different domains, the model can more accurately distinguish similar activities that were previously easily confused. While precision and recall tendencies of some activities may vary, the overall improvement in F1 scores indicates a well-balanced trade-off between precision and recall, resulting in strong performance. The enhanced F1 scores for each activity after domain fusion confirm the model’s improved ability to recognize a variety of activities more accurately. Through these experiments, we verify that using the domain fusion algorithm helps resolve conflicts between labels and enables more precise classification.
@@ VI-E4 Comparison with CNN models
For the comparison with domain fusion of CNN models as shown in Fig. 17(b), we only use the ALERT dataset since RaDA provides only the single domain data. We apply the same CNN models in parallel as feature extractors, passing both range and frequency domain data through each model. We then concatenate the features from each model and pass the concatenated features to the classifier. Here, we use the adjusting factor $\beta$ for the frequency domain, consistent with the domain fusion method of ISA-ViT. Therefore, we also train $\beta$ to achieve the best accuracy in the domain fusion of all CNN models. In a fair comparison, using the best performance of each method, our proposed scheme outperforms the CNN models.
VII Conclusion
@@
This study addresses the critical issue of distracted driving, a significant contributor to traffic accidents worldwide, by proposing a novel Driver Activity Recognition (DAR) framework utilizing IR-UWB radar. IR-UWB technology overcomes the limitations of traditional methods, including privacy concerns, noise sensitivity, and interference issues. Despite its potential, two key challenges are identified: the lack of comprehensive datasets reflecting real-world driving conditions and the difficulty of adapting state-of-the-art models, such as Vision Transformers (ViTs), to process non-image UWB data.
To address these challenges, we developed the ALERT dataset, a large-scale collection of 10,220 samples representing seven diverse distracted driving activities in real-world driving environments. This dataset includes both range and frequency domain data, enabling robust experimentation and setting a new benchmark for realistic DAR studies. Additionally, we introduced ISA-ViT, a tailored model designed to seamlessly process UWB data of varying input sizes by resizing data without loss and adapting pre-trained positional embedding vectors. The integration of domain fusion further enhances performance by combining complementary features from the range and frequency domains.
Through extensive experiments, we analyzed the optimality, trends, and trade-offs of the ALERT dataset parameters across algorithms and demonstrated the effectiveness of the proposed framework. ISA-ViT achieved a classification accuracy of 76.28% and a distracted driving detection accuracy of 97.35%. These results underscore the transformative potential of UWB radar and transformer-based models for advancing DAR research.
By making the ALERT dataset publicly available and providing comprehensive benchmarking results, this study lays a foundation for future research, fostering innovation and collaboration within the DAR community. Ultimately, the proposed methods and dataset aim to contribute to safer road environments by mitigating distracted driving incidents.
[Uncaptioned image] Jeongjun Park received the B.S. degree from the Department of Information Communications Engineering, Hankuk University of Foreign Studies (HUFS), in 2018, and the M.S. and Ph.D. degree in Electrical and Computer Engineering from Seoul National University (SNU), Seoul, South Korea, in 2020 and 2025, respectively. He is currently with System LSI Business in Samsung Electronics. His research interests include the 5G/6G communication technologies, IoT, UWB, and AI.
[Uncaptioned image] Sunwook Hwang received the B.S. degree in Electrical Engineering from Pohang University of Science and Technology (POSTECH) in 2016 and the Ph.D. in Electrical and Computer Engineering from Seoul National University (SNU), Seoul, Korea, in August 2023. He served as a Post-Doctoral Researcher at SNU from September 2023 to August 2024. He is currently with Samsung Electronics, where his work focuses on Neural Processing Unit (NPU) design. His research interests include autonomous driving and computer vision.
[Uncaptioned image] Hyeonho Noh received the B.S. and the M.S. degrees in electrical engineering from Ulsan National Institute of Science and Technology (UNIST), Ulsan, South Korea, in 2018 and 2021, respectively, and received the Ph.D. degree in electrical engineering with the Pohang University of Science and Technology (POSTECH), Pohang, South Korea, in 2024. From Oct. 2024 to Jun. 2025, he worked as a post-doctoral researcher in the Department of Electrical Engineering, Seoul National University (SNU), Seoul, South Korea. He is currently an Assistant Professor with the Department of Information and Communication Engineering, Hanbat National University, Daejeon, South Korea. His research interests include integrated sensing and communication and future wireless systems with deep neural networks.
[Uncaptioned image] Jin Mo Yang received the B.S. degree in electronic engineering from Hanyang University in 2018. He is currently pursuing the Ph.D. degree in electrical and computer engineering with Seoul National University. His research interests include video analytics, edge computing, and test-time adaptation (TTA) for evolving vision domains.
[Uncaptioned image] Hyun Jong Yang is an associate professor in the department of electrical and computer engineering, Seoul National University (SNU), Seoul, Korea. He received the B.S. degree in electrical engineering from Korea Advanced Institute of Science and Technology (KAIST), Korea, in 2004, and the M.S. and Ph.D. degrees in electrical engineering from KAIST, in 2006 and 2010, respectively. From Aug. 2010 to Aug. 2011, he was a research fellow at Korea Research Institute of Ships & Ocean Engineering (KRISO), Korea. From Oct. 2011 to Oct. 2012, he worked as a post-doctoral researcher in the Electrical Engineering Department, Stanford University, Stanford, CA. From Oct. 2012 to Aug. 2013, he was a Staff II Systems Design Engineer, Broadcom Corporation, Sunnyvale, CA, where he developed physical-layer algorithms for LTE-A MIMO receivers. In addition, he was a delegate of Broadcom in 3GPP RAN1 standard meetings. From Sept. 2013 to July 2020, he was an assistant/associate professor in the School of Electrical and Computer Engineering, UNIST, Korea. From July 2020 to Aug. 2024, he was an associate professor in the department of electrical engineering, Pohang University of Science and Technology (POSTECH), Pohang, Korea. Since Sept. 2024, he has been an associate professor in the department of electrical and computer engineering, Seoul National University (SNU), Seoul, Korea. His fields of interests are signal processing, wireless communications, and machine learning. Dr. Yang is an Editor for IEEE Wireless Communications Letters (WCL) and IEEE Internet of Things Journal (IoTJ), an Associate Editor-in-Chief for the Journal of Korean Institute of Communications and Information Sciences, and a Program Track Chair for the IEEE Consumer Communications & Networking Conference (CCNC) 2023 2025.
[Uncaptioned image] Saewoong Bahk (Senior Member, IEEE) received the B.S. and M.S. degrees in electrical engineering from Seoul National University (SNU), in 1984 and 1986, respectively, and the Ph.D. degree from the University of Pennsylvania, in 1991. He was with AT&T Bell Laboratories as a Member of Technical Staff, from 1991 to 1994, where he had worked on network management. From 2009 to 2011, he served as the Director of the Institute of New Media and Communications. He is currently a Professor at SNU. He has been leading many industrial projects on 3G through 6G and IoT connectivity supported by the Korean industry. He has published more than 300 technical articles and holds more than 100 patents. He was a recipient of the KICS Haedong Scholar Award, in 2012. He was President of the Korean Institute of Communications and Information Sciences (KICS). He served as Chief Information Officer (CIO) of SNU. He was General Chair of the IEEE WCNC 2020 (Wireless Communication and Networking Conference), IEEE ICCE 2020 (International Conference on Communications and Electronics), and IEEE DySPAN 2018 (Dynamic Spectrum Access and Networks). He was TPC Chair of the IEEE VTC-Spring 2014, and General Chair of JCCI 2015. He was Director of the Asia–Pacific Region of the IEEE ComSoc. He is an Editor of the IEEE Network Magazine and IEEE Transactions on Vehicular Technology. He was Co-Editor-in-Chief of the Journal of Communications and Networks (JCN), and an editor of Computer Networks Journal and IEEE Tran. on Wireless Communications. He is a member of the National Academy of Engineering of Korea (NAEK).

### v2

@@ VI-B1 Benchmarking of learning algorithms
VI-C Ablation Studies for ALERT dataset
We conduct various experiments to identify the optimal parameters, trends, and trade-offs for each algorithm when utilizing the ALERT dataset. We provide granular ablations beyond resizing/patch selection, including data-size manipulation, few-shot shots, domain usage (range/frequency/fusion), range-Doppler conversion, and fusion strategy (early vs late), as summarized in Figs. 1012.
VI-C1 Analysis for manipulation of data size
As shown in Fig. 10(a), accuracy increases with a longer observation window. Short windows (1–2 s) provide limited temporal context and thus yield lower performance, whereas 5–10 s windows offer sufficient duration to better capture activity patterns. However, larger windows also increase inference latency and computational cost. Therefore, the observation window should be selected by balancing accuracy, system responsiveness, and practical efficiency.
In Fig. 10(b), we compare using all range bins (about 9 m) versus a cropped driver-centric ROI (<<2.7 m) to assess the effect of long-delayed multipath components. While the benefit depends on the model structure, cropping improves performance for all algorithms except ResNet. Although long-delayed multipath signals may occasionally include contextual cues (e.g., passengers or vehicle type), they typically behave as noise for most algorithms; therefore, cropping is generally beneficial.
We also evaluate lower (<<375 MHz), higher (375–750 MHz), and full frequency bands to identify which region best captures motion characteristics for each activity. The optimal choice varies across algorithms, reflecting differences in feature-extraction capability; thus, these results can guide selecting an appropriate model and frequency band.
@@ VI-C2 Analysis for the number of shots
We present the impact of varying the number of shots on a few-shot adaptation. We adopt few-shot adaptation to reduce the domain gap when transferring the PEV capability learned from large-scale image pre-training to the UWB domain. To obtain a few shot samples, the driver has to capture the activities, which can be burdensome; however, this occurs only once when the system is initially operated. For the few-shot adaptation phase, we utilize various number of shot samples and 10 epochs adaptation, which are enough to fit the model to a driver. During adaptation, we fine-tune the entire model (pre-trained PEVs, backbone, and classifier) for 10 epochs using Adam with the same learning rate as training (model-dependent, 0.00001–0.0001) and an adaptation batch size of 5. As shown in Fig. 10(c), the accuracy increases when all models’ number of shots increases. In our experiments, the few-shot samples are taken only from the Relax behavior, and they are strictly excluded from the testing set to prevent any data leakage. We observe that even with the addition of only about 5 shots, the accuracy gain is significant. In particular, when applying 30 shots to ISA-ViT, it exhibits a maximum performance of 91.75%.
@@ VI-C3 Analysis for the input data domain
In particular, CNN-based models tend to benefit from frequency-domain inputs that provide richer local patterns, whereas transformer-based models (including ISA-ViT) more effectively exploit range-domain spatial cues; domain fusion is most beneficial when the model can balance complementary information from both domains.
@@ VI-C4 Analysis for the range-Doppler data domain
As shown in Fig. 2, ALERT provides UWB representations in the range–time and frequency–time domains and allows users to flexibly reformat the data for different needs. To demonstrate this flexibility, we additionally evaluate the range–Doppler representation, which is used in datasets such as RaDA and is also supported by ALERT. This conversion can be performed using the provided script (ALERT_makeDataset.py) in our public repository (https://github.com/ALERTdataset/ALERT.git).
As shown in Fig. 12, CNN- and RNN-based methods with range–Doppler inputs do not outperform their domain-fusion counterparts in Fig. 9(a). In contrast, ViT-based models (DeiT and ViT) achieve higher performance with range–Doppler inputs than with domain fusion.
This trend is mainly related to input resizing and resolution loss. CNNs and RNNs can process UWB inputs directly without resizing, whereas DeiT and ViT require resizing to $224\times 224$, which reduces resolution. In domain fusion, the time resolution is reduced by more than half for both range–time and frequency–time inputs; converting to range–Doppler also reduces the Doppler resolution by about half. Our results suggest that, for these ViT baselines, losing time resolution is more detrimental than losing Doppler resolution, making range–Doppler relatively preferable to domain fusion.
The proposed ISA-ViT builds on a ViT backbone but introduces a resizing strategy that preserves the original information, enabling improved performance even with range–Doppler inputs. Moreover, ISA-ViT employs complementary fusion of range and frequency features; as shown in Fig. 9(a), this yields higher accuracy than using range–Doppler alone.
@@ VI-D Data Verification
To further verify that the collected signals are not trivially similar across participants even within the same behavior, we quantify inter-participant variability using a prototype-based distance analysis. For each behavior and each participant, we construct a participant-level prototype by averaging all available samples after applying sample-wise normalization to mitigate amplitude/offset differences. We then compute pairwise dissimilarity between participant ($\mathbf{p}$) prototypes using cosine distance, $d_{i,j}=1-\cos(\mathbf{p}_{i},\mathbf{p}_{j})$, which emphasizes similarity in spatiotemporal patterns rather than absolute magnitude. As shown in Fig. 14, the resulting inter-participant distance heatmaps (one per behavior) exhibit diverse off-diagonal values across participant pairs, indicating noticeable participant-dependent variations and supporting that the ALERT dataset is not composed of near-duplicate signals. This complements the t-SNE feature visualization by providing a direct, quantitative view of cross-participant variability at the signal level.
(a) Drive
(b) Relax
(c) Nod
(d) Drink
(e) Phone
(f) Smoke
(g) Panel
Figure 14: Inter-participant cosine-distance heatmaps for seven activities (Drive, Relax, Nod, Drink, Phone, Smoke, and Panel).
@@ VI-E1 Overall performance
We refer to the classification accuracy (76.28%), F1 score in Table VIII, and confusion matrix in Fig. 15(b) to analyze the model’s accuracy from both safety and service perspectives. The classification accuracy reflects the model’s ability to categorize all activities within our proposed framework. The precision and recall of each activity indicate the safety and service factors. If the recall for Drive is low, the system may trigger false alarms, potentially interrupting the driving experience and reducing overall service quality. While false alarms may be frustrating from a service perspective, they do not pose significant safety risks. However, the precision for Drive is directly linked to safety. Low precision could indicate that other potentially dangerous activities, such as using a smartphone, smoking, nodding are incorrectly classified as driving. Thus, we should refer to classification accuracy and F1 score together to analyze the accuracy in perspectives of safety and service.
@@
TABLE VIII: Precision, recall, and F1 score for each label
@@ VI-E2 Impact on resizing method
In the following ISA-ViT-specific evaluations (Figs. 1618, we further analyze the contributions of information-preserving resizing and complementary domain fusion under fair comparison settings. To evaluate ISA-ViT, we conduct experiments with several resizing methods for the ALERT and RaDA datasets. We have already verified that maintaining the $\times14$ sequence of pre-trained PEVs yields better performance than manipulating them in Section V. Therefore, we only compare methods that maintain pre-trained PEVs.
@@ VI-E3 Impact on domain fusion
TABLE IX: Comparison of precision, recall, and F1 score before/after domain fusion
@@
Moreover, Fig. 17 illustrates the accuracy impact of the trained adjusting factor $\beta$ for frequency features. Before training the adjusting factor $\beta$, we achieve a performance of 74.02% with range and frequency features at an equal ratio. However, after jointly training $\beta$ and applying it to domain fusion, we observe an accuracy of 76.28%. This confirms that it is crucial not merely to combine range and frequency features, but to balance them appropriately for optimal fusion. This effect is also reflected at the activity level (Table. IX), where domain fusion improves the F1 score across all activities (e.g., Panel: +13.69, Drink: +10.50, Relax: +9.82, Nod: +9.23), indicating that complementary domain information is particularly helpful for distinguishing visually similar behaviors.
The detailed performance enhancements are presented in Table IX, which shows the precision, recall, and F1 score for each activity before and after applying domain fusion. For all activities, the F1 scores improve after domain fusion is applied. These improvements demonstrate that by combining information from different domains, the model can more accurately distinguish similar activities that were previously easily confused. While precision and recall tendencies of some activities may vary, the overall improvement in F1 scores indicates a well-balanced trade-off between precision and recall, resulting in strong performance. The enhanced F1 scores for each activity after domain fusion confirm the model’s improved ability to recognize a variety of activities more accurately. Through these experiments, we verify that using the domain fusion algorithm helps resolve conflicts between labels and enables more precise classification.
@@ VI-E4 Comparison with CNN models
TABLE X: FLOPs of learning algorithms
Algorithms GoogLeNet ResNet DenseNet MobileNet RNN DeiT ViT ISA-ViT
FLOPs 3.82 G 6.43 G 3.82 G 0.35 G 7.53 G 1.08 G 59.70 G 60.01 G
@@
For the comparison with domain fusion of CNN models as shown in Fig. 18(b), we only use the ALERT dataset since RaDA provides only the single domain data. We apply the same CNN models in parallel as feature extractors, passing both range and frequency domain data through each model. We then concatenate the features from each model and pass the concatenated features to the classifier. Here, we use the adjusting factor $\beta$ for the frequency domain, consistent with the domain fusion method of ISA-ViT. Therefore, we also train $\beta$ to achieve the best accuracy in the domain fusion of all CNN models. In a fair comparison, using the best performance of each method, our proposed scheme outperforms the CNN models. Taken together, these studies provide component-level justification for ISA-ViT by showing that information-preserving resizing with preserved pre-trained PEVs and complementary domain fusion with a learned $\beta$ are key to robust performance across UWB input shapes.
VII Discussion
VII-A Points of Attention for ALERT Dataset
The ALERT dataset is the first UWB-based dataset that comprehensively captures distracted driving activities in a real-driving environment, and we expect it to support future DAR research. Nevertheless, we acknowledge several limitations.
First, ALERT was collected using a single vehicle (a mid-sized sedan), as acquiring additional vehicle types (e.g., SUVs, compact cars, and large vehicles) was logistically challenging. To reduce potential bias from cabin geometry, we fixed the sensor mounting configuration and analyzed a driver-centric ROI by cropping range bins around the driver, which mitigates vehicle-dependent reflections. Moreover, as shown in Sec. VI, long-delayed multipath components that often carry environmental information tend to act as noise for most algorithms, suggesting that region of interest (ROI) cropping can further suppress environment/vehicle variations. While multi-vehicle experiments are left for future work, the radar-to-driver setup was kept consistent and the participant cohort still provides meaningful variability, so we expect the main findings to remain valid.
Second, only four of the nine participants collected data for all labels. This was mainly due to the smoking label: to avoid unrealistic samples, we excluded non-smokers mimicking smoking, and thus not all participants could contribute to that class. Nonetheless, each label was collected from a majority of participants, which we consider sufficient for model training and evaluation.
VII-B Implications and Further Explorations of ISA-ViT
ISA-ViTachieves substantially higher performance than other models with only a small additional cost over standard ViT, which is a key advantage. Built on the ViT architecture to transfer its representational power to UWB domains, ISA-ViT inevitably requires more computation than CNN/RNN baselines. While this trade-off is acceptable given the accuracy gain, low-latency response remains important in DAR scenarios.
Table X summarizes the computational cost of each model. ISA-ViT improves accuracy by 22.68% over ViT with only a 0.3 GFLOPs increase; however, real-time inference may still be challenging on resource-limited edge devices.
In future work, we will reduce computation via pruning, quantization, and knowledge distillation, and explore efficient transformer variants tailored for UWB data to retain high accuracy with lower cost for edge deployment.
VIII Conclusion
@@
This study addresses distracted driving—a major cause of traffic accidents—by proposing an IR-UWB-radar-based Driver Activity Recognition (DAR) framework. IR-UWB mitigates key limitations of conventional approaches, such as privacy concerns and susceptibility to noise and interference. However, its adoption has been hindered by (i) the lack of comprehensive datasets collected under real-driving conditions and (ii) the difficulty of adapting state-of-the-art models, such as Vision Transformers (ViTs), to non-image UWB data.
To overcome these challenges, we develop the ALERT dataset, comprising 10,220 samples of seven distracted driving activities recorded in real-driving environments. The dataset provides both range- and frequency-domain representations, enabling robust evaluations and realistic benchmarking. We also propose ISA-ViT, which supports variable-size UWB inputs by resizing the data without loss and adapting pre-trained positional embeddings. Furthermore, domain fusion improves recognition by combining complementary features from the range and frequency domains.
Extensive experiments investigate dataset-parameter effects across algorithms and validate the proposed framework. ISA-ViT achieves 76.28% classification accuracy and 97.35% distracted-driving detection accuracy, demonstrating the promise of UWB radar and transformer-based modeling for DAR.
By publicly releasing ALERT with comprehensive benchmarks, this work establishes a foundation for future DAR research and aims to support safer driving by reducing distracted driving incidents.